# Jr. AI Scientist and Its Risk Report: Autonomous Scientific Exploration from a Baseline Paper

### v2

@@ 3.3.2 Stage1: Idea Implemention
The system manages four experimental nodes running in parallel, each responsible for implementing and testing a proposed idea independently. Within each node, the coding agent receives the baseline code and a research idea, and writes a directly executable script named proposed_method.py. Once the coding agent finishes writing the implementation, the system sequentially runs proposed_method.py and plot.py. If a result file is successfully generated, the codebase is marked as Non-Buggy; otherwise, it is labeled as Buggy If a visualization image is also produced, it is further marked as Non-Plot-Buggy; otherwise, as Plot-Buggy. Each iteration of this process—coding and execution—is counted as one trial and is repeated until a bug-free implementation is obtained.
As shown in Figure˜2, if any node completes successfully without encountering bugs, its codebase is carried forward to Stage 2. If all nodes fail, the system selects the next nodes, either initializing new nodes from the baseline code or debugging previously generated buggy codebases. When debugging buggy codebases, we provide the coding agent with detailed runtime feedback, such as standard output and error messages, to guide iterative debugging until the issue is resolved. We set this stage to run for 12 iterations.
@@ 3.4 Writing Phase
We primarily used a coding agent (e.g. Claude Code (claude_code)) as a writing agent for the writing process. As shown in Figure˜3, the Writing Phase consists of three stages—Draft Writing, Reflection, and Adjustment. Below, we describe the resources provided to the writing agent and the details of each stage.
@@ 3.4.2 Draft Writing
Rationale of Writing Orders. Following AI Scientist v2 (yamada2025ai), we initially generated the entire paper at once, but this resulted in a decline in the quality of the Method section. We therefore adopted a step-by-step writing process to improve overall consistency and quality. When determining the writing order, we considered it most important for the writing agent to first accurately understand and describe the proposed method, as this understanding serves as the foundation for correctly writing other sections. Therefore, we instruct the writing agent to focus exclusively on accurately writing the Method section. In addition, for the paper structure, we followed the approach of AI-Researcher (tang2025ai) and introduced an intermediate step of summarizing the paper structure, in which the writing agent briefly outlines the content of each section before full-paper generation. These refinements made it possible to produce more consistent and accurate descriptions throughout a paper.
Collection of BibTeX Entries. To ensure accurate citation, it is essential to collect a complete and correct set of BibTeX entries. Following AI Scientist v1 (lu2024ai), we use the Semantic Scholar API to retrieve BibTeX records. However, this approach alone often yields an insufficient number of references. To address this limitation, we adopt a practical strategy commonly used by human researchers: using the baseline paper’s BibTeX file as a starting point. This approach allows the system to gather a sufficient number of references while also can expect correct citation by referring to the baseline’s LaTeX source. Since the baseline’s BibTeX file does not include the entry for the baseline reference set, we explicitly add it to the reference set.
@@ 3.4.3 Reflection
(2) Reflection on Formatting and Presentation. Following AI Scientist v2 (yamada2025ai), we also introduce a reflection phase focused on formatting and presentation quality. In this phase, the writing agent generates feedback such as: “Are there any LaTeX syntax errors or style violations we can fix? Refer to the chktex output below”, or “Are there short sections (one or two sentences) that could be combined into a single paragraph?” This process helps produce a final draft that is well-formatted and stylistically consistent.
(3) Feedback Generation and Reflection on Figures. Following AI Scientist v2 (yamada2025ai), we perform figure-level reflection in the refinement stages by integrating a Large Multimodal Model (LMM)-based feedback mechanism. This process aims to improve the quality, clarity, and alignment of generated figures, captions, and their corresponding textual interpretations. Specifically, the LMM is used to identify figures that are uninformative or make little contribution to the paper’s scientific value, and such figures are either removed or moved to the Appendix. This ensures that all figures presented in the main paper contain adequate informational value. To achieve this, we provide the LMM with the paper’s abstract, figure captions, and figure images to generate targeted feedback, which is then used to guide the reflection and revision process.
(4) Feedback Generation and Reflection from AI Reviews. Following CycleResearcher (weng2024cycleresearcher), we adopt a review-based reflection, where the system improves its manuscript based on reviewer feedback. In this step, the writing agent revises the generated paper according to reviewer comments such as “the Method section is unclear”, “important parameter details are missing”, or “the writing is overly verbose”. For generating such feedback, we employ AI reviewers in AI Scientist v1 (text-only evaluation) and v2 (evaluation including figures). These AI reviewers use GPT-4o (gpt4o) and are prompted to evaluate papers in the official NeurIPS review format.
@@ 3.4.4 Adjustment: Page-length Adjustment
We also introduced a new design to the page-length adjustment process. In AI Scientist v2 (yamada2025ai), when the generated paper exceeded the predefined page limit, the system attempted to adjust the length within a single LLM call. However, this approach often resulted in over-trimming, causing the paper to become significantly shorter than the target length. To address this issue, our method performs iterative and gradual page-length reduction until the manuscript reaches the target length, thereby improving the stability of page adjustment. As a result, the final papers consistently fell within 1 page of the specified page limit. We set the page layout to 8 pages.
@@ 4.1 Experimental Setting
Baseline Paper Selection. In selecting the baseline papers, we were concerned about unpredictable impact and computational cost. For the former, we considered the potential societal impact of accelerating research through AI Scientist systems, which might pose a risk of confusing the research fields. To mitigate such risks, we prioritized our own papers and, for other works, selected only those for which we obtained explicit permission from the original authors. For the latter, we selected papers that require relatively few GPU hours, ensuring that the experiments can be conducted even in our academic laboratories. Although this limits the ability to conduct large-scale experiments, it does not undermine our objectives to evaluate the capability and risks of AI Scientist systems during their development.
As a result, we selected three papers: LoCoOp (NeurIPS2023) (miyai2023locoop) and GL-MCM (IJCV2025) (miyai2025gl) in the field of OOD detection, and Min-K%++ (ICLR2025 spotlight) (zhang2024min) in the field of pre-training data detection for LLMs. LoCoOp is a few-shot learning method with CLIP (radford2021learning), GL-MCM is an inference-only method with CLIP, and Min-K%++ is an inference-only task with LLMs. Both research areas have recently attracted increasing attention (shi2023detecting, miyai2024generalized). Accurately evaluating how much current AI Scientist systems can advance these research fields is crucial for deepening the understanding of their capabilities and limitations.
Human Involvement. In our framework, humans were involved only in verifying the outputs. Publicly available papers are often curated and therefore may not accurately represent the typical quality of each system’s outputs. Therefore, for evaluation, following this common practice, we selected the papers that appeared to be of the highest quality among several generated ones. We include three generated papers in the Appendix.
@@ 4.2 Results with Public AI Reviewers
Comparison Methods. As comparison methods, we used papers generated by existing AI Scientist systems. Specifically, we included AI Scientist-v1 (lu2024ai), AI Scientist-v2 (yamada2025ai), AI-Researcher (tang2025ai), CycleResearcher (weng2024cycleresearcher), and Zochi (zochi2025).
Evaluation Metrics. For evaluation, we employed DeepReviewer (zhu2025deepreview), an AI model designed to comprehensively assess research papers in a manner similar to expert reviewers. DeepReviewer not only performs summarization and scoring, but also reproduces an internal reasoning chain that sequentially examines aspects such as novelty, reliability, and clarity. In its novelty verification stage, it leverages external tools such as the Semantic Scholar API and OpenScholar (asai2024openscholar) to perform evidence-based literature search and information retrieval, allowing for evidence-driven evaluations. Importantly, our method does not incorporate DeepReviewer into the reflection process, ensuring that the obtained scores are unbiased.
Results. We present our experimental results in Table˜2(c). The table shows the average scores across multiple papers (overall score), the score of the paper with the highest rating, and that of the paper with the lowest rating. From these results, we observe that our papers outperform the publicly available AI Scientist papers in all criteria (Soundness, Presentation, Contribution, and Rating).
@@ 5 Agents4Science Conference Submission
Agents4Science (agents4science) is a conference jointly organized by Stanford University and Together AI, where AI systems serve as both the primary authors and reviewers of research papers. The first edition of the conference was held in 2025. It is the first venue in which AI authorship is not only allowed but required, enabling open evaluation of AI-generated research and the development of guidelines for responsible AI participation in science. This conference targets a wide range of AI-driven contributions, including papers authored by AI Scientists, as well as those that allow human involvement. The conference provides an ideal platform for evaluating our work, so we submitted our paper to this venue to receive feedback from its AI reviewers.
@@
The AI reviewers used in Agents4Science are based on GPT-5 (gpt5), Gemini 2.5 (gemini2.5pro), and Claude Sonnet 4 (anthropic2025claude4). They tune these models through in-context learning using review samples from ICLR 2024 and ICLR 2025.
@@
While the performance improvement is not necessarily limited, it is indeed true that the Jr. AI Scientist still has considerable room for further enhancement. To achieve higher performance, it would be necessary to increase the number of experimental trials and explore more innovative search strategies for selecting experimental nodes.
@@ 6 Authors Evaluation
We conducted an internal review of the generated papers with authors. Recent AI-generated papers include some degree of manual post-editing (zochi2025, weng2025deepscientist), and only a few studies have carefully examined the raw, unedited outputs of AI systems (yamada2025ai). However, evaluating the raw, unedited outputs is essential for accurately understanding the current limitations of AI Scientist systems.
The review here does not evaluate whether the paper has the level of contribution, impact, or experimental results required for acceptance at a conference. Instead, our review focuses on whether the writing contains misinterpretations of results, incorrect methodological descriptions, inaccurate citations, or hallucinations. Therefore, we cross-check the manuscripts against the actual code and experimental results to precisely identify such issues. The issues of our review are summarized as follows. More detailed reviews for each paper are provided in the Appendix.
@@ 7.1 Idea Generation
From this perspective, more extensive validation was conducted in some recent works (liu2025alphago, weng2025deepscientist). For example, the concurrent work DeepScientist (weng2025deepscientist) performed a comprehensive large-scale study. They report that, out of approximately 5,000 unique scientific ideas generated, only 21 ultimately led to genuine scientific innovations. Our experiments require less time because our limitation analysis is effective, and our goal is modest, aiming to find one successful idea rather than exploring more successful ideas.
@@ 7.2 Experiment
Because the coding agent is unaware of domain-specific conventions, it often improves performance in undesirable or invalid ways. This issue frequently appeared in the experiments on GL-MCM (miyai2025gl), which we describe in detail below.
Background. GL-MCM is a task of zero-shot out-of-distribution (OOD) detection (miyai2024generalized). OOD detection aims to distinguish between samples belonging to a predefined in-distribution (ID) class set (e.g., the 1000 classes of ImageNet) and those belonging to classes with different semantics (yang2024generalized). In the GL-MCM setting, the model uses CLIP (radford2021learning) and is required to discriminate between ID and OOD data without any training, given only the ID class names.
@@ 7.4 Review
Current AI Reviewers primarily evaluate the written content of papers and lack any mechanism to detect discrepancies between the text and the actual experimental results. For instance, even if all the reported ablation studies were fabricated, there is no way for the reviewer to identify such inconsistencies. A similar observation was also reported in (jiang2025badscientist). To address this issue, it would be necessary to develop a reviewing agent that can access and analyze all associated code files and result data in addition to the manuscript. Developing AI reviewers that can incorporate not only textual information but also experimental code and data will be an important direction for future research.
@@ 8 Conclusion
Acknowledgements
We would like to thank Jingyang Zhang for permission to use Min-K%++. This work was partially supported by JSPS KAKENHI 25H01164.
@@ Acknowledgements
The appendix includes the three papers generated in this study:

### v4

@@ 3.3.2 Stage1: Idea Implemention
The system manages four experimental nodes running in parallel, each responsible for implementing and testing a proposed idea independently. Within each node, the coding agent receives the baseline code and a research idea, and writes a directly executable script named proposed_method.py. Once the coding agent finishes writing the implementation, the system sequentially runs proposed_method.py and plot.py. If a result file is successfully generated, the codebase is marked as Non-Buggy; otherwise, it is labeled as Buggy If a visualization image is also produced, it is further marked as Non-Plot-Buggy; otherwise, as Plot-Buggy. Each iteration of this process, coding and execution, is counted as one trial and is repeated until a bug-free implementation is obtained.
As shown in Figure˜2, if any node completes successfully without encountering bugs, its codebase is carried forward to Stage 2. During execution, four experimental nodes are selected. If all currently running nodes fail, the system selects the next nodes, either by initializing new nodes from the baseline code or by debugging previously generated buggy codebases. When debugging buggy codebases, we provide the coding agent with detailed runtime feedback, such as standard output and error messages, to guide iterative debugging until the issue is resolved. We set this stage to run for 12 iterations.
@@ 3.4 Writing Phase
We primarily used a coding agent (e.g. Claude Code (Anthropic, 2025a)) as a writing agent for the writing process. As shown in Figure˜3, the Writing Phase consists of three stages—Draft Writing, Reflection, and Adjustment. Below, we describe the resources provided to the writing agent and the details of each stage.
@@ 3.4.1 Preliminary: Resources Provided to Writing Agent
The only materials that humans need to prepare for each paper generation are the LaTeX source files and the baseline codebase of the baseline paper. The CLAUDE.md file and the conference LaTeX template are shared across all experiments. The code and experimental files for the proposed methods are automatically generated during the experimental stage. Using these inputs, the AI agent automatically writes all sections and captions in the paper.
@@ 3.4.2 Draft Writing
Rationale of Writing Orders. Following AI Scientist v2 (Yamada et al., 2025), we initially generated the entire paper at once, but this resulted in a decline in the quality of the Method section. We therefore adopted a step-by-step writing process to improve overall consistency and quality. When determining the writing order, we considered it most important for the writing agent to first accurately understand and describe the proposed method, as this understanding serves as the foundation for correctly writing other sections. Therefore, we instruct the writing agent to focus exclusively on accurately writing the Method section. In addition, for the paper structure, we followed the approach of AI-Researcher (Tang et al., 2025) and introduced an intermediate step of summarizing the paper structure, in which the writing agent briefly outlines the content of each section before full-paper generation. These refinements made it possible to produce more consistent and accurate descriptions throughout a paper.
Collection of BibTeX Entries. To ensure accurate citation, it is essential to collect a complete and correct set of BibTeX entries. Following AI Scientist v1 (Lu et al., 2024), we use the Semantic Scholar API to retrieve BibTeX records. However, this approach alone often yields an insufficient number of references. To address this limitation, we adopt a practical strategy commonly used by human researchers: using the baseline paper’s BibTeX file as a starting point. This approach allows the system to gather a sufficient number of references while also can expect correct citation by referring to the baseline’s LaTeX source. Since the baseline’s BibTeX file does not include the entry for the baseline reference set, we explicitly add it to the reference set.
@@ 3.4.3 Reflection
(2) Reflection on Formatting and Presentation. Following AI Scientist v2 (Yamada et al., 2025), we also introduce a reflection phase focused on formatting and presentation quality. In this phase, the writing agent generates feedback such as: “Are there any LaTeX syntax errors or style violations we can fix? Refer to the chktex output below”, or “Are there short sections (one or two sentences) that could be combined into a single paragraph?” This process helps produce a final draft that is well-formatted and stylistically consistent.
(3) Feedback Generation and Reflection on Figures. Following AI Scientist v2 (Yamada et al., 2025), we perform figure-level reflection in the refinement stages by integrating a Large Multimodal Model (LMM)-based feedback mechanism. This process aims to improve the quality, clarity, and alignment of generated figures, captions, and their corresponding textual interpretations. Specifically, the LMM is used to identify figures that are uninformative or make little contribution to the paper’s scientific value, and such figures are either removed or moved to the Appendix. This ensures that all figures presented in the main paper contain adequate informational value. To achieve this, we provide the LMM with the paper’s abstract, figure captions, and figure images to generate targeted feedback, which is then used to guide the reflection and revision process.
(4) Feedback Generation and Reflection from AI Reviews. Following CycleResearcher (Weng et al., 2025a), we adopt a review-based reflection, where the system improves its manuscript based on reviewer feedback. In this step, the writing agent revises the generated paper according to reviewer comments such as “the Method section is unclear”, “important parameter details are missing”, or “the writing is overly verbose”. For generating such feedback, we employ AI reviewers in AI Scientist v1 (text-only evaluation) and v2 (evaluation including figures). These AI reviewers use GPT-4o (Hurst et al., 2024) and are prompted to evaluate papers in the official NeurIPS review format.
@@ 3.4.4 Adjustment: Page-length Adjustment
We also introduced a new design to the page-length adjustment process. In AI Scientist v2 (Yamada et al., 2025), when the generated paper exceeded the predefined page limit, the system attempted to adjust the length within a single LLM call. However, this approach often resulted in over-trimming, causing the paper to become significantly shorter than the target length. To address this issue, our method performs iterative and gradual page-length reduction until the manuscript reaches the target length, thereby improving the stability of page adjustment. As a result, the final papers consistently fell within 1 page of the specified page limit. We set the page layout to 8 pages.
@@ 4.1 Experimental Setting
Baseline Paper Selection. In selecting the baseline papers, we were concerned about unpredictable impact and computational cost. For the former, we considered the potential societal impact of accelerating research through AI Scientist systems, which might pose a risk of confusing the research fields. To mitigate such risks, we selected only those for which we obtained explicit permission from the original authors. For the latter, we selected papers that require relatively few GPU hours, ensuring that the experiments can be conducted even in our academic laboratories. Although this limits the ability to conduct large-scale experiments, it does not undermine our objectives to evaluate the capability and risks of AI Scientist systems during their development.
As a result, we selected three papers: LoCoOp (NeurIPS2023) (Miyai et al., 2023) and GL-MCM (IJCV2025) (Miyai et al., 2025b) in the field of out-of-distribution (OOD) detection (a task that aims to detect semantic classes outside the predefined set of semantic classes), and Min-K%++ (ICLR2025 spotlight) (Zhang et al., 2025) in the field of pre-training data detection for LLMs. LoCoOp is a few-shot learning method with CLIP (Radford et al., 2021), GL-MCM is an inference-only method with CLIP, and Min-K%++ is an inference-only task with LLMs. Both research areas have recently attracted increasing attention (Shi et al., 2024; Miyai et al., 2025a). Accurately evaluating how much current AI Scientist systems can advance these research fields is crucial for deepening the understanding of their capabilities and limitations.
Cost. Claude Code is the most resource-intensive agent in our pipeline. However, by using the Claude Code Max plan (USD 200 per month), it is possible to run two experiments in parallel within the usage limit, enabling paper generation without incurring substantial cost.
Human Involvement. In our framework, humans were involved only in verifying the outputs. Publicly available papers are often curated and therefore may not accurately represent the typical quality of each system’s outputs. Therefore, for evaluation, following this common practice, we selected the papers that appeared to be of the highest quality among six generated ones. As the selection criterion, one of the authors who is highly familiar with the baseline paper carefully examined the content of the generated papers, the obtained results, and the code, and selected those that were considered to be of high quality. We include three generated papers in the supplementary material.
@@ 4.2 Results with Public AI Reviewers
Comparison Methods. As comparison methods, we used papers generated by existing AI Scientist systems. Specifically, we included AI Scientist-v1 (Lu et al., 2024), AI Scientist-v2 (Yamada et al., 2025), AI-Researcher (Tang et al., 2025), CycleResearcher (Weng et al., 2025a), and Zochi (Intology, 2025).
Evaluation Metrics. For evaluation, we employed DeepReviewer (Zhu et al., 2025a), an AI model designed to comprehensively assess research papers in a manner similar to expert reviewers. DeepReviewer is a 14B-parameter language model built by fine-tuning Phi-4 on the DeepReview-13K dataset, which contains structured, human-like review reasoning trajectories. During evaluation, DeepReviewer focuses on technical soundness, experimental validity, and logical consistency, grounding its judgments in explicit evidence from the manuscript. This design enables efficient and reliable evaluation while maintaining strong alignment with human reviewer judgments. This adopts standardized scoring, where the overall rating score is from 1 to 10, and soundness, presentation, and contribution are from 1 to 4. Compared to existing AI reviewers (Weng et al., 2025a; Lu et al., 2024; Yamada et al., 2025), DeepReviewer exhibits substantially stronger alignment with human reviewer evaluations, enabling efficient and reliable assessment of AI-generated paper reviews. For the implementation, we utilize a single A100 80G GPU.
Results. We evaluated papers generated by our system (an extension paper of LoCoOp is shown in Figure˜4). All generated papers, along with detailed author comments, are provided in the supplementary material. We present our experimental results in Table˜2(c). Here, among the three papers we generated, the LoCoOp extension paper received a score of 6, the GL-MCM extension paper received a score of 5, and the Min-K%++ extension paper received a score of 6.25. The table shows the average scores across multiple papers (overall score), the score of the paper with the highest rating, and that of the paper with the lowest rating. From these results, we observe that our papers outperform the publicly available AI Scientist papers in all criteria (Soundness, Presentation, Contribution, and Rating).
Score Distribution among Generated Papers. Figure˜5 shows the score distribution among all our generated papers before the author selection. First, the average score over all 18 papers is 5.30. Although this is lower than the score of 5.75 reported in Table˜2(c), it can still be considered higher than those of existing AI-generated papers. Second, while the authors’ selection criteria show a similar tendency to the review results produced by DeepReviewer, the selected papers are not necessarily those with the highest review scores. For example, in the case of LoCoOp, the selected paper has a score of 6, even though papers with higher scores of 6.25 and 6.50 exist. Upon manually reviewing these higher-scoring papers, we found that they contain more hallucinations in terms of numerical values and claims (Detailed hallucinations and risks are discussed in Section˜6 and Section˜7. ). These findings indicate that evaluation based solely on review scores is insufficient, and highlight the importance of author inspection and honest reporting of hallucinations and associated risks.
@@ 5 Agents4Science Conference Submission
Agents4Science (Zou et al., 2025; Bianchi et al., 2025) is a conference jointly organized by Stanford University and Together AI, where AI systems serve as both the primary authors and reviewers of research papers. The first edition of the conference was held in 2025. It is the first venue in which AI authorship is not only allowed but required, enabling open evaluation of AI-generated research and the development of guidelines for responsible AI participation in science. This conference targets a wide range of AI-driven contributions, including papers authored by AI Scientists, as well as those that allow human involvement. The conference provides an ideal platform for evaluating our work, so we submitted our paper to this venue to receive feedback from its AI reviewers.
@@
This comment is valid. While Jr. AI Scientist achieved higher scores than the baseline, the performance gap is not significant enough to claim a substantial improvement. To address this limitation, it would be necessary to increase the number of experimental trials and explore more innovative search strategies for selecting experimental nodes.
@@ 6 Authors Evaluation
We conducted an internal review of the generated papers with authors. Recent AI-generated papers include some degree of manual post-editing (Intology, 2025; Weng et al., 2025b), and only a few studies have carefully examined the raw, unedited outputs of AI systems (Yamada et al., 2025). However, evaluating the raw, unedited outputs is essential for accurately understanding the current limitations of AI Scientist systems.
The review here does not evaluate whether the paper has the level of contribution, impact, or experimental results required for acceptance at a conference. Instead, our review focuses on whether the writing contains misinterpretations of results, incorrect methodological descriptions, inaccurate citations, or hallucinations. Therefore, we cross-check the manuscripts against the actual code and experimental results to precisely identify such issues. The issues of our review are summarized as follows. More detailed reviews for each paper are provided in the supplementary material.
@@ 7.1 Idea Generation
From this perspective, more extensive validation was conducted in some recent works (Liu et al., 2025; Weng et al., 2025b). For example, the concurrent work DeepScientist (Weng et al., 2025b) performed a comprehensive large-scale study. They report that, out of approximately 5,000 unique scientific ideas generated, only 21 ultimately led to genuine scientific innovations. Our experiments require less time because our limitation analysis is effective, and our goal is modest, aiming to find one successful idea rather than exploring more successful ideas.
@@ 7.2 Experiment
Because the coding agent is unaware of domain-specific conventions, it often improves performance in undesirable or invalid ways. This issue frequently appeared in the experiments on GL-MCM (Miyai et al., 2025b), which we describe in detail below.
Background. GL-MCM is a task of zero-shot out-of-distribution (OOD) detection (Miyai et al., 2025a). OOD detection aims to distinguish between samples belonging to a predefined in-distribution (ID) class set (e.g., the 1000 classes of ImageNet) and those belonging to classes with different semantics (Yang et al., 2024). In the GL-MCM setting, the model uses CLIP (Radford et al., 2021) and is required to discriminate between ID and OOD data without any training, given only the ID class names.
@@ 7.4 Review
Current AI Reviewers primarily evaluate the written content of papers and lack any mechanism to detect discrepancies between the text and the actual experimental results. For instance, even if all the reported ablation studies were fabricated, there is no way for the reviewer to identify such inconsistencies. A similar observation was also reported in (Jiang et al., 2025). To address this issue, it would be necessary to develop a reviewing agent that can access and analyze all associated code files and result data in addition to the manuscript. Developing AI reviewers that can incorporate not only textual information but also experimental code and data will be an important direction for future research.
@@ 8 Conclusion
Limitations and Future Work
Use of Multiple Agents. We built our AI Scientist on top of a single coding agent (i.e. Claude Code) and a single family of large language models. We did not explore the potential benefits of combining multiple LLMs or coding agents to further improve performance. Investigating multi-agent or multi-model configurations remains an important direction for future work.
Fragility of Novelty Verification via Semantic Scholar. Our novelty verification relies on Semantic Scholar, which remains fragile and incomplete. How to more reliably ensure and validate the novelty of generated research is an important open problem and a key challenge for future research.
Addressing Observed Risks and Identified Limitations. In this study, we place emphasis on accurately reporting observed risks and limitations identified through projects. Developing methods to mitigate these risks and address the identified limitations will be an important direction for future efforts.
Broader Impact Statement
Through this work, we developed a Jr. AI Scientist. However, we do not recommend the use of Jr. AI Scientist or similar systems in the preparation of actual conference submission papers. The primary objective of this paper is to help the community gain a clearer and more comprehensive understanding of the current status of AI Scientists. As discussed throughout the paper, AI Scientists entail various risks. Therefore, when preparing manuscripts for conference submission, it is essential that all content be carefully inspected and validated by human authors prior to publication.
Furthermore, we argue that researchers working on AI Scientists bear a responsibility to report risks, including concrete failure cases and negative outcomes. The study of AI Scientists has significant implications for the academic ecosystem. As these systems continue to advance in capability and autonomy, it is essential that developers remain fully aware of the potential risks they may introduce and commit to their responsible development and deployment. Responsible advancement requires not only the reporting of successful outcomes, but also the systematic disclosure of limitations, failure cases, and potential systemic risks.
This paper adheres to such a responsible research principle. In addition to presenting the capabilities of AI Scientists, we systematically document the risks and shortcomings observed in our study. Through this transparent reporting, we hope to help the community gain a clearer and more comprehensive understanding of the current status of AI Scientists.
@@
Appendix A Public Release of Codebase and Detail Prompts
We plan to gradually release a codebase. This decision is motivated by the potential risks associated with making the Jr. AI Scientist easy to use. As discussed in the main paper, AI Scientist systems, including the Jr. AI Scientist, also involve various risks, and the potential impact on the academic community must be carefully considered. Therefore, we plan to release a codebase after evaluating the risks that this system may pose to the research community.
Appendix B Main Prompts for Jr. AI Scientist
An example of the prompt in the Idea Generation Phase is shown below. Since the Idea Generation Phase consists of three stages, different prompts are provided for each stage.
B.1 Idea Generation
B.2 Experimentation
An example of the prompt given to the Coding Agent to make it write the code required for the Experiment Phase is shown below. Since the Experiment Phase consists of four stages, different prompts are provided for each stage. In addition, even within the same stage, elements such as <memory> may vary slightly across different executions of the agent.
The <introduction>, <directory_structure>, and <research_idea> elements are shared across all stages.
The <research_idea> directly contains the research idea generated in the Idea Generation Phase, providing the Coding Agent with a high-level contextual overview of the functionality to be implemented.
In Stage 2 and Stages 3–4, <performance_improvement_idea> and <ablation_study_idea> are inserted, respectively. These guide the implementation strategy.
The <memory> field embeds a summary of the experiments conducted so far within the corresponding stage. This includes results and metrics from successful experiments, as well as causes of failed experiments and error logs, summarized by the LLM.
The <required_functionality> specifies the main functionalities that the Coding Agent is expected to implement. While this content differs across stages, the differences are limited to a short descriptive text and the name of the entry-point file. In Stage 1, where the proposed method is implemented, this field is set to <proposed_method_implementation>. In Stage 2, which focuses on improving the performance of the proposed method, the required functionality is specified as <improved_proposed_method>.
B.3 Writing
The prompt used for Draft Writing in the Writing Phase is shown below. The portion denoted as {RESEARCH IDEA} is identical to the text presented in B.2. {RESOURCE OVERVIEW} is the resource explanation provided to the writing agent. {PLOT LIST} and {PLOT DESCRIPTION} correspond to experimental results and may vary across different executions.
Next, we present the prompt used for Reflection.
Appendix C Detail of Agents4Science Reviews
This section presents detailed reviews from Agents4Science. In Section˜5, we focused on reviews that address the fundamental limitations of the Jr. AI Scientist. In contrast, this section provides detailed reviews to clarify what is required for AI Scientist–generated papers to be accepted at future Agents4Science conferences. (If the goal is just acceptance, including human intervention is the fastest way.) As noted in Section˜5, the papers submitted were generated by an earlier version of the Jr. AI Scientist and are therefore not identical to the paper introduced in this work.
C.1 LoCoOp Extention Paper
Reviewer 1
Reviewer 2
Reviewer 3
C.2 Min-K%++ Extention Paper
Reviewer 1
Reviewer 2
Reviewer 3
C.3 GL-MCM Extention Paper
Reviewer 1
Reviewer 2
Reviewer 3
Appendix D Generated Papers
The section includes the three papers generated in this study: